### v1

@@ 2.2.2 X-ray Polarimetry
X-ray polarimetry represents an independent tool for measuring the geometry of the innermost accretion flow, the X-ray corona, and the black hole spin. While for the coronal emission the X-ray polarization is expected to be sensitive to the geometry and motion of the corona [83, 84, 85, 86, 87, 88], for the purpose of measuring the black hole spin the X-ray polarization of emission from the thin disc is more relevant as the relativistic signatures are imprinted in the level and energy dependence of the polarization degree and angle [89, 90, 91]. In the case of strongly polarized accretion discs seen at high (almost edge-on) inclination, the polarization signal of returning radiation is expected to be significant [91, 92, 93, 94].
@@
To demonstrate the capability of eXTP in constraining the black hole spin by X-ray polarimetry, we perform simulations of eXTP/PFA polarisation observations of a bright black hole X-ray binary in the thermal state. For the simulations, we assume a black hole X-ray binary with a flux of 1  Crab in 2–8 keV and at a distance of 3 kpc. The mass of the black hole is $7M_{\odot}$. The observer is located at an inclination of $60^{\circ}$. The energy spectra and polarization is computed with the kynbbrr model [89, 92, 93, 94] and a disk albedo of 1 is assumed. The results are shown in Fig. 7 (polarization degree and angle in left and right panels, respectively). By looking at the simulated data for an extremely rotating black hole ($a_{*}=0.998$; green crosses) and a non-rotating ($a_{*}=0$; blue crosses) black hole, as well as the model predictions for different values of black hole spins, it is evident that eXTP observations will be able to put stringent constraints on black hole spin.
Figure 7: The simulated eXTP/PFA observations for polarization degree (PD) and polarization angle (PA) for extremely rotating (green crosses, $a=0.998$) and non-rotating (blue crosses, $a=0$) black holes. The simulations of 500 ks observations of a generic 1 Crab black hole X-ray binary (in 2-8 keV) have been prepared with the ixpeobssim v31.0.3 updated for the eXTP/PFA responses. Black hole mass of 7 $M_{\odot}$, inclination of $60^{\circ}$, distance of 3 kpc and high disk albedo of 1 was assumed. The model predictions with Stokes parameters binned in 22 energy bins for different values of the black hole spin is also shown. Both the simulated observations and the model predictions were created with plot polfrac and plot polang commands in XSPEC v12.14.1d.
@@ 2.2.3 Thermal continuum fitting
The X-ray thermal continuum fitting method offers an effective way to measure black hole spin by analyzing thermal emission in the X-ray band from a multi-temperature accretion disk surrounding accreting black holes. The temperature distribution in the accretion disk and various gravitational effects are highly sensitive to the mass and spin of the central black hole, allowing the spectral energy distribution to provide sufficient details to constrain the black hole’s properties. This approach requires precise measurements of the black hole mass $M$, the inclination angle of the system $i$, and the distance $D$. To date, the spins of more than 20 black holes have been successfully measured using this method [66, 67].
@@
The black hole spectra for eXTP/SFA, eXTP/pnCCD, and XTI/NICER were simulated using the relativistic continuum model kerrbb, assuming a 300 ks exposure time. A total of 40 simulated spectra were generated, covering the full range of black hole spin from 0.000 to 0.998. A disk inclination of $i=57^{\circ}$, a source distance of $D=7.47$ kpc, and a black hole mass of $M=8.46M_{\odot}$ were adopted. The inclination corresponds to the expectation value ($\bar{i}=1$ radian or $57^{\circ}$) for randomly oriented orbital planes, computed from the probability distribution $P(i)\propto\sin i$ on a sphere. The mass and distance were chosen to correspond to the median values in the observed distribution of black hole X-ray binaries [97]. The mass accretion rate $M_{dd}$ was adjusted to yield a luminosity corresponding to 5% of the Eddington limit. By default, a canonical spectral hardening factor of 1.7 was assumed. To account for self-irradiation and limb-darkening effects, both corresponding flags were set to 1. The power-law photon index was set to 2.0, and the power-law normalization was tuned to match an unabsorbed flux of $1.2\times 10^{-8}$ erg cm -2 s-1 in the energy band 2-10 keV. A hydrogen column density of $nH=1.0\times 10^{22}\ cm^{-2}$ was applied. The energy range used in recovery is 2.0-10.0 keV for all instruments. No systematic uncertainties were added to the simulated spectra for any of the instruments.
@@
To estimate confidence intervals for the freely fitted parameters ($nH$, $a_{*}$, $M_{dd}$, $\Gamma$, and $Norm_{PL}$), Markov Chain Monte Carlo (MCMC) sampling was performed using the Goodman-Weare algorithm. Each chain was sampled for $10^{6}$ steps, and convergence was evaluated using the $\hat{R}$ diagnostic. All model parameters were successfully recovered from the simulated spectra in all configurations. For each spin value investigated, the eXTP / SFA configuration yielded the tightest constraints, with confidence intervals reduced by a factor of 2 to 3 compared to those obtained with NICER/XTI. While the pnCCD produced slightly broader intervals than NICER/XTI, both remained within the same order of magnitude, indicating overall comparable performance. Representative results are shown in Figure 8 for spin values of 0.000 and 0.998, with red, green, and blue denoting the confidence intervals for eXTP/SFA, eXTP/pnCCD, and NICER/XTI, respectively.
@@
Figure 8: MCMC sampling results for simulated spectra with spin assumed at extreme values of 0.000 and 0.998. The simulation paramters were set as $M=8.46M_{\odot}$, $D=7.47$ kpc, $i=57.0^{\circ}$, while other parameters were adjusted to achieve a 5% Eddington luminosity and an unabsorbed flux of $1.2\times 10^{-8}$ erg cm-2 s-1 in the 2–10 keV energy range. No systematic uncertainties were added. Confidence contours are presented at the 68% and 95% confidence levels, with red, green and blue denoting the confidence intervals for eXTP/SFA, eXTP/pnCCD, and NICER/XTI, respectively. The tightest constraints were obtained for eXTP/SFA.
@@
Near the surface of the black hole, the gravitational field is extremely intense. The final region of spacetime, known as the plunging region, is where infalling matter cannot be halted and inevitably descends into the black hole. By analyzing three-dimensional general relativistic magnetohydrodynamic (GRMHD) simulations of black hole accretion disks and measuring black hole spin using the continuum spectrum fitting method, both approaches yield consistent results. This suggests that the current use of the Novikov-Thorne thin disk model does not significantly constrain the accuracy of spin measurements obtained through continuous spectral fitting methods [98]. It has been found that for black holes with high spin, the combined emission from the plunging region can produce a weak power-law tail in the high-energy spectrum, which may be the origin of corona emission [99]. Through the study of the outburst state of the black hole X-ray binary MAXI J0637-430, it was found that an additional thermal component is required to fit the spectrum. This component can be self-consistently explained by photons emitted from the accretion flow within the innermost stable circular orbit of the black hole [100].
@@ 2.2.4 X-ray Quasi Periodic Oscillations
X-ray Quasi Periodic Oscillations (QPOs) have been routinely observed in stellar mass black holes at low (0.1–10  Hz) and high (few hundred Hz) frequencies, within 10-20% of the relativistic precession, orbital, and epicyclic frequencies in the inner disk. These frequencies correspond to the fundamental frequencies of orbital motion in strong-field general relativity (GR). The low and high-frequency QPOs have sometimes been observed together, in combinations of frequencies consistent with those expected from the multiple relativistic signals associated. These observations suggest that QPOs may be indicative of oscillations occurring within specific radii, thereby offering a potential means to accurately probe relativistic dynamic of particles in the vicinity of a black hole event horizon. Furthermore, QPOs can provide measurements of black hole spin, a crucial parameter that influences the observed frequencies.
@@
The Relativistic Precession Model (RPM) has been pivotal in linking QPOs to black hole parameters. It explains black hole QPOs as motions of accreting matter in the strong gravitational field of a Kerr black hole, matching observed QPO frequencies to general-relativistic orbital time-scales. In this model, a low-frequency QPO (typically a few Hz) is produced by Lense–Thirring frame dragging that causes the nodal precession of matter around the black hole, while high-frequency QPOs (tens to hundreds of Hz) arise from the Keplerian orbital frequency and the relativistic periastron precession of matter. Such types of motion are assumed to occur at the same characteristic radius, often associated with the inner truncation radius of the geometrically thin accretion disk. These fundamental frequencies depend on the black hole’s mass and spin, so detecting a “triplet” of QPOs (one low-frequency plus two high-frequency signals) allows one to infer the black hole mass and dimensionless spin by solving the GR frequency relations [103, 104, 105].
@@
The leading model for low-frequency QPOs assumes that the corona (also sometimes referred to as the hot flow) undergoes nodal (Lense-Thirring) precession around the black hole spin axis due to the relativistic frame dragging effect, whilst the disk remains stationary due to viscous forces [106]. eXTP is ideally suited to two distinctive tests of this model.
@@
First, the model predicts that the iron line should rock between Doppler red shifted and Doppler blue shifted as the corona preferentially illuminates different disk azimuths throughout its precession cycle [107]. Fig 9a shows an example for two QPO phases. The disk images (top and middle) demonstrate the illumination profile at the two phases, where the rainbow color scheme is used for disk patches illuminated by a flux above a threshold value. The corresponding line profiles for the same two phases are shown in the bottom panel. For axi-symmetric illumination, the line profile would instead be independent of QPO phase. Since the QPO frequency varies stochastically, it is very difficult to phase-fold. Instead, a QPO phase-dependence in spectral shape is optimally inferred in the Fourier domain [108], by measuring the amplitude and phase lag with respect to a reference band as a function of photon energy for the first and second QPO harmonics. Fig 9b demonstrates that a phase-dependence of the iron line profile creates distinctive features in the iron line region of the amplitude and lag spectra.
@@
The high throughput and spectral resolution of the SFA will enable the first firm detections of asymmetric disk illumination. To demonstrate this, we simulate a 50 ks eXTP observation of a 250 mCrab black hole X-ray binary with a QPO at 0.4 Hz with quality factor $Q=8$. We assume that the photon index $\Gamma$ and the reflection fraction $f_{R}$ oscillate with QPO phase around their mean values of $\Gamma=1.8$ and $f_{R}=0.4$, and we model asymmetric illumination with the asymmetry parameters $A_{1}$ and $A_{2}$ (using the best fitting values for H 1743-322 [110]). We also set $\log\xi=3$, $r_{\rm in}=30\leavevmode\nobreak\ r_{g}$, $i=70^{\circ}$ for the disk ionization parameter, inner radius and inclination angle, respectively. We simulate using the cross spectral error formulae of ref. [111]. Fig 9b shows the resulting amplitude and lag spectra for the first two QPO harmonics. The features in the $\sim 5-8$ keV range indicative of a QPO phase-dependent iron line profile can be clearly seen in the simulated data. The dashed gray lines represent the best fitting null hypothesis model with $A_{1}$ and $A_{2}$ set to zero. We see that this model can clearly be ruled out by the synthetic eXTP observation. Fig 9c shows the statistical confidence with which the asymmetry parameters can be constrained by fitting the input model back to the simulated data [^2] We fit to the real and imaginary parts of the cross spectrum instead of directly to the amplitude and phase lag, which is statistically favorable.. The cross represents the best fitting values and the contours represent $68.27\%$, $95.45\%$ and $99.73\%$ confidence. The null hypothesis of axi-symmetric illumination ($A_{1}=A_{2}=0$) is ruled out with $\gg 5\leavevmode\nobreak\ \sigma$ significance.
@@
Figure 9: a) Disk images representing asymmetric illumination by a precessing corona at two precession phases (top and middle) and the corresponding iron line profiles (bottom). The QPO phase-dependent illumination of the disk is parameterized by an analytic function with asymmetry parameters that we set to the best fitting values for H 1743-322 [110] of $A_{1}=0.93$ and $A_{2}=3.50$. b) Simulated 50 ks eXTP observation of a 250 mCrab black hole X-ray binary with a QPO at 0.4 Hz. We show the amplitude and phase lag of the first and second QPO harmonics as a function of energy (as labeled). The simulated data are unfolded around the best fitting model (red stepped lines). The gray dashed lines represent the null-hypothesis model. c) Statistical confidence contours (red: $68.27\%$, green: $95.45\%$ and blue: $99.73\%$) on the measurement of the asymmetry parameters from fitting the input model back to the simulated data.
@@
The second test, enabled by the PFA, is to search for modulations of the polarization degree (PD) and polarization angle (PA) on the QPO period. Such polarization QPOs are expected from a precessing corona in analogy with the rotating vector model for radio pulsars. eXTP holds many key advantages over IXPE concerning this test. One is the larger effective area, not just of the PFA, but also of the SFA, which can be utilized by cross correlating the PFA signal with the SFA signal [112]. Another is the superior telemetry. Telemetry limitations restrict each IXPE observation to a maximum of $\sim 15$ million counts, which will not be the case for eXTP. eXTP will also be better able to slew quickly to outbursting black hole XRBs, enabling it to observe when the QPO properties are optimal, and the higher orbit will enable longer uninterrupted observations.
@@
Detection of a polarization QPO requires a long exposure of a bright source, therefore we simulate a 150 ks eXTP observation of a 300 mCrab black hole X-ray binary in the hard state exhibiting a QPO with centroid frequency $0.4$ Hz and quality factor $Q=8$. We calculate the QPO phase dependence of flux, PD and PA assuming that the corona is a torus of outer radius $20\leavevmode\nobreak\ r_{g}$, misaligned with the disk by an angle $\beta=10^{\circ}$, which is viewed from inclination $i=70^{\circ}$ (i.e. Fig 6 of [113]). We assume that the time-averaged PD is $4\%$. We simulate using the Fourier detection method of [112], using an improved treatment of statistical errors [111]. This method consists of extracting light curves selected for different modulation angle bins and, for each of these light curves, measuring the QPO amplitude and phase lag with respect to a reference light curve. This is in direct analogy to spectral-timing techniques, except the different time series considered are selected on modulation angle rather than on energy channel. We simulate for 30 modulation angle bins, and we use the SFA for the reference band. We simulate $10^{4}$ realizations of the same observation. Fig 10 shows the results of one representative realization of the simulation. The points are the simulated eXTP data, the gray lines represent the null hypothesis of constant PD and PA, and the red lines the best fitting model of variable PD and PA. We use an F-test for each realization of the simulation to find that the best fitting model is preferred over the null-hypothesis with a statistical confidence of $5.2\leavevmode\nobreak\ \sigma$. We calculate this significance from the median F-statistic of all $10^{4}$ realizations.
2.3 Testing General Relativity
Figure 11: Constraints on the black hole spin parameter $a_{*}$ and the Johannsen deformation parameter $\alpha_{13}$ ($\alpha_{13}=0$ corresponds to the Kerr solution of General Relativity and $\alpha_{13}\neq 0$ implies deviations from General Relativity) from a simulated 30 ks NICER observation (in blue), a simulated 30 ks eXTP observation (in red), and a simulated 300 ks eXTP observation (in green) of a bright black hole X-ray binary. In the simulations, we assume that the source is a Kerr black hole with spin parameter $a_{*}=0.5$ and the inclination angle of the disk is 70 deg. The other parameters have the values used in the simulations in Fig. 5. For every measurement, we show the 68%, 90%, and 99% confidence level curves for two parameters.
The spacetime metric around an astrophysical black hole is normally thought to be approximated well by the Kerr solution. However, there are models beyond General Relativity allowing for macroscopic deviations from the Kerr geometry. Testing the Kerr nature of astrophysical black holes is thus a crucial test of General Relativity in the strong field regime [60].

### v2

@@ 2.2.2 X-ray Polarimetry
X-ray polarimetry represents an independent tool for measuring the geometry of the innermost accretion flow, the X-ray corona, and the black hole spin. While for the coronal emission the X-ray polarization is expected to be sensitive to the geometry and motion of the corona [96, 97, 98, 99, 100, 101], for the purpose of measuring the black hole spin the X-ray polarization of emission from the thin disc is more relevant as the relativistic signatures are imprinted in the level and energy dependence of the polarization degree and angle [102, 103, 104]. In the case of strongly polarized accretion disks seen at high (almost edge-on) inclination, the polarization signal of returning blackbody radiation from the disk is expected to be significant [104, 105, 106, 107].
@@
To demonstrate the capability of eXTP in constraining the black hole spin by X-ray polarimetry, we perform simulations of eXTP/PFA polarisation observations of a bright black hole X-ray binary in the soft (thermal-dominated) state. For the simulations, we assume a black hole X-ray binary with a flux of 1  Crab in 2–8 keV and at a distance of 3 kpc. The mass of the black hole is $7M_{\odot}$. The observer is located at an inclination of $60^{\circ}$. The energy spectra and polarization is computed with the kynbbrr model [102, 105, 106, 107] and a disk albedo of 1 is assumed. The results are shown in Fig. 7 (polarization degree and angle in left and right panels, respectively). By looking at the simulated data for an extremely rotating black hole ($a_{*}=0.998$; green crosses) and a non-rotating ($a_{*}=0$; blue crosses) black hole, as well as the model predictions for different values of black hole spins, it is evident that eXTP observations will be able to put stringent constraints on black hole spin.
Figure 7: The simulated eXTP/PFA observations for polarization degree (PD) and polarization angle (PA) for extremely rotating (green crosses, $a_{*}=0.998$) and non-rotating (blue crosses, $a_{*}=0$) black holes. The simulations of 500 ks observations of a generic 1 Crab black hole X-ray binary (in 2-8 keV) in the soft state, have been prepared with the ixpeobssim v31.0.3 updated for the eXTP/PFA responses. Black hole mass of 7 $M_{\odot}$, inclination of $60^{\circ}$, distance of 3 kpc and high disk albedo of 1 was assumed. The model predictions with Stokes parameters binned in 22 energy bins for different values of the black hole spin is also shown. Both the simulated observations and the model predictions were created with plot polfrac and plot polang commands in XSPEC v12.14.1d.
@@ 2.2.3 Thermal continuum fitting
The X-ray thermal continuum fitting method offers an effective way to measure black hole spin by analyzing thermal emission in the X-ray band from a multi-temperature accretion disk surrounding accreting black holes. This method requires the source to be in the high-soft state where the thin-disk model assumptions hold, with disk emission contributing dominantly (>>75% [46]) to the total flux and weak scattering/reflection components. The temperature distribution in the accretion disk and various gravitational effects are highly sensitive to the mass and spin of the central black hole, allowing the spectral energy distribution to provide sufficient details to constrain the black hole’s properties. This approach requires precise measurements of the black hole mass $M$, the inclination angle of the system $i$, and the distance $D$. To date, the spins of more than 20 black holes have been successfully measured using this method [78, 79].
@@
The black hole spectra for three configurations: eXTP/SFA(full array of 5 SDD modules + 1 pnCCD module), eXTP/SFA-I(one pnCCD only), and NICER/XTI were simulated using the relativistic continuum model kerrbb, assuming a 300 ks exposure time. A total of 40 simulated spectra were generated, covering the full range of black hole spin from 0.000 to 0.998. A disk inclination of $i=57^{\circ}$, a source distance of $D=7.47$ kpc, and a black hole mass of $M=8.46M_{\odot}$ were adopted. The inclination corresponds to the expectation value ($\bar{i}=1$ radian or $57^{\circ}$) for randomly oriented orbital planes, computed from the probability distribution $P(i)\propto\sin i$ on a sphere. The mass and distance were chosen to correspond to the median values in the observed distribution of black hole X-ray binaries [110]. The mass accretion rate $M_{\rm{dd}}$ was adjusted to yield a luminosity corresponding to 5% of the Eddington limit. By default, a canonical spectral hardening factor of 1.7 was assumed. To account for self-irradiation and limb-darkening effects, both corresponding flags in kerrbb were set to 1. The self-irradiation, i.e. due to returning blackbody radiation, is assumed to be fully-absorbed (i.e. albedo 0) and thus reprocessed at the local blackbody temperature. This is inconsistent with the assumed fully reflecting disk in the polarization modelling of the soft state shown in the previous section. Modelling of combined SFA and PFA data from eXTP will require the development and application of fully self-consistent spectro-polarimetric models, including more realistic disk atmospheres. The power-law photon index was set to 2.0, and the power-law normalization was set to ensure that the disk emission contribution excced 75% of the total unabsorbed flux in the 2-10 keV energy band. In the actual synthetic spectra, the disk component contributed 80-85% of the total unabsorbed flux. The absorbed flux ranging from $3.4\times 10^{-9}$ erg cm -2 s-1 to $8.7\times 10^{-9}$ erg cm -2 s-1 after accouting for the specified column density of $N_{\rm{H}}=1.0\times 10^{22}cm^{-2}$. No systematic uncertainties were added to the simulated spectra for any of the instruments.
@@
To estimate confidence intervals for the freely fitted parameters ($N_{\rm{H}}$, $a_{*}$, $M_{\rm{dd}}$, $\Gamma$, and $Norm_{\rm{PL}}$), Markov Chain Monte Carlo (MCMC) sampling was performed using the Goodman-Weare algorithm. Each chain was sampled for $10^{6}$ steps, and convergence was evaluated using the $\hat{R}$ diagnostic. All model parameters were successfully recovered from the simulated spectra in all configurations. For each spin value investigated, the eXTP / SFA configuration yielded the tightest constraints, with confidence intervals reduced by a factor of 3 to 4 compared to those obtained with NICER/XTI. While the SFA-I produced slightly broader intervals than NICER/XTI, both remained within the same order of magnitude, indicating overall comparable performance. Representative results are shown in Fig. 8 for spin values of 0.000 and 0.998, with red, green, and blue denoting the confidence intervals for eXTP/SFA, eXTP/SFA-I, and NICER/XTI, respectively.
@@
Figure 8: MCMC sampling results for simulated spectra with spin assumed at extreme values of 0.000 and 0.998. The simulation paramters were set as $M=8.46M_{\odot}$, $D=7.47$ kpc, $i=57.0^{\circ}$, while other parameters were adjusted to achieve a 5% Eddington luminosity and the disk component contributing 80–85% of the unabsorbed flux. The flux of the brightest spectrum corresponds to 0.07 Crab in the 2-10 keV energy band. No systematic uncertainties were added. Confidence contours are presented at the 68% and 95% confidence levels, with red, green and blue denoting the confidence intervals for eXTP/SFA, eXTP/SFA-I, and NICER/XTI, respectively. The tightest constraints were obtained for eXTP/SFA.
@@
Near the surface of the black hole, the gravitational field is extremely intense. The final region of spacetime, known as the plunging region, is where infalling matter cannot be halted and inevitably descends into the black hole. It has been found that for black holes with high spin, the combined emission from the plunging region can produce a weak power-law tail in the high-energy spectrum, which may be the origin of corona emission [111]. Fitting synthetic spectra from three-dimensional general relativistic magnetohydrodynamic (GRMHD) simulations with current relativistic thermal emission models (e.g., kerrbb for thin disks and slimbh for slim disks) could help to understand systematics in spin measurements arising from the neglected plunging region emission and other modeling simplifications. Studies of this kind have not converged to a consistent understanding yet, with some simulations shown negligible systematic errors [112] while others indicate a significant overestimation of the black hole spin [113, 114]. Semi-analytical models that include plunging region emission have been developed [115], and their application to data has shown that black hole spin measurements can be lower than those obtained using standard models in some cases (e.g., MAXI J0637-430 [116], MAXI J1820+070 [115] and M33 X-7 [117]).
Another aspect concerning the reliability of standard disk models is the stability. While the standard disk models predict the disk to be viscously and thermally unstable in the radiation pressure-dominated regime [118], observations of BHXRBs in the corresponding soft state show very stable X-ray emission. A possible solution to this discrepancy is the inclusion of magnetic fields [119], which could alter the disk structure and temperature profile [113, 120].
@@ 2.2.4 X-ray Quasi Periodic Oscillations
X-ray Quasi Periodic Oscillations (QPOs) have been routinely observed in stellar mass black holes at low (0.1–10  Hz) and high (few hundred Hz) frequencies, within 10-20% of the relativistic precession, orbital, and epicyclic frequencies in the inner disk. These frequencies correspond to the fundamental frequencies of orbital motion in strong-field GR. The low and high-frequency QPOs have sometimes been observed together, in combinations of frequencies consistent with those expected from the multiple relativistic signals associated. These observations suggest that QPOs may be indicative of oscillations occurring within specific radii, thereby offering a potential means to accurately probe relativistic dynamic of particles in the vicinity of a black hole event horizon. Furthermore, QPOs can provide measurements of black hole spin, a crucial parameter that influences the observed frequencies.
@@
The Relativistic Precession Model (RPM) has been pivotal in linking QPOs to black hole parameters. It explains black hole QPOs as motions of accreting matter in the strong gravitational field of a Kerr black hole, matching observed QPO frequencies to general-relativistic orbital time-scales. In this model, a low-frequency QPO (typically a few Hz) is produced by Lense–Thirring frame dragging that causes the nodal precession of matter around the black hole, while high-frequency QPOs (tens to hundreds of Hz) arise from the Keplerian orbital frequency and the relativistic periastron precession of matter. Such types of motion are assumed to occur at the same characteristic radius, often associated with the inner truncation radius of the geometrically thin accretion disk. These fundamental frequencies depend on the black hole’s mass and spin, so detecting a “triplet” of QPOs (one low-frequency plus two high-frequency signals) allows one to infer the black hole mass and dimensionless spin by solving the GR frequency relations [123, 124, 125, 126].
@@
The leading model for low-frequency QPOs assumes that the corona corresponds to an inner hot accretion flow, with orbital motion misaligned with the black hole spin. The misalignment causes the corona to undergo solid-body nodal (Lense-Thirring) precession around the black hole spin axis due to the relativistic frame dragging effect, whilst the disk remains stationary due to viscous forces [127]. eXTP is ideally suited to two distinctive tests of this model.
@@
First, the model predicts that the iron line should rock between Doppler red shifted and Doppler blue shifted as the corona preferentially illuminates different disk azimuths throughout its precession cycle [128]. Fig. 9a shows an example for two QPO phases. The disk images (top and middle) demonstrate the illumination profile at the two phases, where the rainbow color scheme is used for disk patches illuminated by a flux above a threshold value. The corresponding line profiles for the same two phases are shown in the bottom panel. For axi-symmetric illumination, the line profile would instead be independent of QPO phase. Since the QPO frequency varies stochastically, it is very difficult to phase-fold. Instead, a QPO phase-dependence in spectral shape is optimally inferred in the Fourier domain [129], by measuring the amplitude and phase lag with respect to a reference band as a function of photon energy for the first and second QPO harmonics. Fig. 9b demonstrates that a phase-dependence of the iron line profile creates distinctive features in the iron line region of the amplitude and lag spectra.
@@
The high throughput and spectral resolution of the SFA will enable the first firm detections of asymmetric disk illumination. To demonstrate this, we simulate a 50 ks eXTP observation of a 250 mCrab black hole X-ray binary with a QPO at 0.4 Hz with quality factor $Q=8$. We assume that the photon index $\Gamma$ and the reflection fraction $f_{R}$ oscillate with QPO phase around their mean values of $\Gamma=1.8$ and $f_{R}=0.4$, and we model asymmetric illumination with the asymmetry parameters $A_{1}$ and $A_{2}$ (using the best fitting values for H 1743-322 [131]). We also set $\log\xi=3$, $r_{\rm in}=30\leavevmode\nobreak\ R_{g}$, $i=70^{\circ}$ for the disk ionization parameter, inner radius and inclination angle, respectively. We simulate using the cross spectral error formulae of ref. [132]. Fig. 9b shows the resulting amplitude and lag spectra for the first two QPO harmonics. The features in the $\sim 5-8$ keV range indicative of a QPO phase-dependent iron line profile can be clearly seen in the simulated data. The dashed gray lines represent the best fitting null hypothesis model with $A_{1}$ and $A_{2}$ set to zero. We see that this model can clearly be ruled out by the synthetic eXTP observation. Fig. 9c shows the statistical confidence with which the asymmetry parameters can be constrained by fitting the input model back to the simulated data [^5] We fit to the real and imaginary parts of the cross spectrum instead of directly to the amplitude and phase lag, which is statistically favorable.. The cross represents the best fitting values and the contours represent $68.27\%$, $95.45\%$ and $99.73\%$ confidence. The null hypothesis of axi-symmetric illumination ($A_{1}=A_{2}=0$) is ruled out with $\gg 5\leavevmode\nobreak\ \sigma$ significance.
@@
Figure 9: (a) Disk images representing asymmetric illumination by a precessing corona at two precession phases (top and middle) and the corresponding iron line profiles (bottom). The QPO phase-dependent illumination of the disk is parameterized by an analytic function with asymmetry parameters that we set to the best fitting values for H 1743-322 [131] of $A_{1}=0.93$ and $A_{2}=3.50$. (b) Simulated 50 ks eXTP observation of a 250 mCrab black hole X-ray binary with a QPO at 0.4 Hz. We show the amplitude and phase lag of the first and second QPO harmonics as a function of energy (as labeled). The simulated data are unfolded around the best fitting model (red stepped lines). The gray dashed lines represent the null-hypothesis model. (c) Statistical confidence contours (red: $68.27\%$, green: $95.45\%$ and blue: $99.73\%$) on the measurement of the asymmetry parameters from fitting the input model back to the simulated data.
@@
The second test, enabled by the PFA, is to search for modulations of the polarization degree (PD) and polarization angle (PA) on the QPO period. Such polarization QPOs are expected from a precessing corona in analogy with the rotating vector model for radio pulsars. eXTP holds many key advantages over IXPE concerning this test. One is the larger effective area, not just of the PFA, but also of the SFA, which can be utilized by cross correlating the PFA signal with the SFA signal [133]. Another is the superior telemetry. Telemetry limitations restrict each IXPE observation to a maximum of $\sim 15$ million counts, which will not be the case for eXTP. eXTP will also be better able to slew quickly to outbursting BHXRBs, enabling it to observe when the QPO properties are optimal, and the higher orbit will enable longer uninterrupted observations.
Detection of a polarization QPO requires a long exposure of a bright source, therefore we simulate a 150 ks eXTP observation of a 300 mCrab black hole X-ray binary in the hard state exhibiting a QPO with centroid frequency $0.4$ Hz and quality factor $Q=8$. We calculate the QPO phase dependence of flux, PD and PA assuming that the corona is a torus of outer radius $20\leavevmode\nobreak\ R_{g}$, misaligned with the disk by an angle $\beta=10^{\circ}$, which is viewed from inclination $i=70^{\circ}$ (i.e. Fig 6 of [134]). We assume that the time-averaged PD is $4\%$. We simulate using the Fourier detection method of [133], using an improved treatment of statistical errors [132]. This method consists of extracting light curves selected for different modulation angle bins and, for each of these light curves, measuring the QPO amplitude and phase lag with respect to a reference light curve. This is in direct analogy to spectral-timing techniques, except the different time series considered are selected on modulation angle rather than on energy channel. We simulate for 30 modulation angle bins, and we use the SFA for the reference band. We simulate $10^{4}$ realizations of the same observation. Fig. 10 shows the results of one representative realization of the simulation. The points are the simulated eXTP data, the gray lines represent the null hypothesis of constant PD and PA, and the red lines the best fitting model of variable PD and PA. We use an F-test for each realization of the simulation to find that the best fitting model is preferred over the null-hypothesis with a statistical confidence of $5.2\leavevmode\nobreak\ \sigma$. We calculate this significance from the median F-statistic of all $10^{4}$ realizations.
@@
2.3 Testing GR
Figure 11: Constraints on the black hole spin parameter $a_{*}$ and the Johannsen deformation parameter $\alpha_{13}$ ($\alpha_{13}=0$ corresponds to the Kerr solution of GR and $\alpha_{13}\neq 0$ implies deviations from GR) from a simulated 30 ks NICER observation (in blue), a simulated 30 ks eXTP observation (in red), and a simulated 300 ks eXTP observation (in green) of a bright black hole X-ray binary. In the simulations, we assume that the source is a Kerr black hole with spin parameter $a_{*}=0.5$ and the inclination angle of the disk is 70 deg. The other parameters have the values used in the simulations in Fig. 5. For every measurement, we show the 68%, 90%, and 99% confidence level curves for two parameters.
The spacetime metric around an astrophysical black hole is normally thought to be approximated well by the Kerr solution. However, there are models beyond GR allowing for macroscopic deviations from the Kerr geometry. Testing the Kerr nature of astrophysical black holes is thus a crucial test of GR in the strong field regime [72].